OpenStudy (antaniece):

What is the value of 5^3 i^9?

OpenStudy (jjcoolja12):

this is what i got

OpenStudy (mathstudent55):

What is 5^3 ?

OpenStudy (antaniece):

really lol i have no idea how to do it

OpenStudy (jjcoolja12):

(53)(i9) =125i

OpenStudy (antaniece):

125

OpenStudy (jjcoolja12):

that is what i got

OpenStudy (mathstudent55):

Have you learned exponents? What does an exponent of 3 tell you to do?

OpenStudy (mathstudent55):

Correct. 5^3 = 5 * 5 * 5 = 125

OpenStudy (mathstudent55):

Now we need to deal with i^9

OpenStudy (antaniece):

im just confused on the i^9

OpenStudy (mathstudent55):

I'll explain i^9 and i to any whole number power. Any number raised to the 0 power (except 0) equals 1. That means i^0 = 1 By definition, a number raised to the 1 power equals itself, so i^1 = i By definition, i^2 = -1 Ok so far?

OpenStudy (antaniece):

yeah

OpenStudy (mathstudent55):

So far we have this: \(i^0 = 1\) \(i^1 = i\) \(i^2 = -1\) Now we need i^3. \(i^3 = i^2 \times i = -1 \times i = -i\) Now we have \(i^3 = -i\) Now we do one more, \(i^4\) \(i^4 = i^3 \times i = -i \times i = -i^2 = -(-1) = 1\) Now let's write the entire thing again, so we can see a pattern: \(i^0 = 1\) \(i^1 = i\) \(i^2 = -1\) \(i^3 = -i\) \(i^4 = 1\)

OpenStudy (mathstudent55):

Once we see that \(i^4\) is the same as \(i^0\), we can tell that \(i^5\) will be the same as \(i^1\), etc. This pattern will continue. Base Exponent Result i multiple of 4 1 i (multiple of 4) + 1 i i (multiple of 4) + 2 -1 i (multiple of 4) + 3 -i

OpenStudy (mathstudent55):

Now we compare 9 to a multiple of 4. 8 is a multiple of 4, since 8 = 2 * 4 9 = 8 + 1 9 is a (multiple of 4) plus 1, in better English, 9 is 1 more than a multiple of 4. For an exponent of 1 more than a multiple of 4, i raised to that exponent equals i \(\Large i^9 = i\)

OpenStudy (mathstudent55):

Example: What is \(\Large i^{253}\) ? Answer: 253/4 = 63.25, so 253 = 4 * 63 + 1 Since 253 is 1 more than a multiple of 4, \(\Large i^{253} = i\)